# The Coralscapes Dataset: Semantic Scene Understanding in Coral Reefs

### v1

@@ VI Conclusion and Future Work
We present the Coralscapes dataset for semantic understanding in reef scenes, which is the first general purpose dataset of coral scenes that are densely and consistently annotated by trained experts. Coralscapes is representative of the challenges of semantic segmentation in reef scenes, including complex scenes with challenging environmental conditions. By benchmarking state-of-the-art computer vision models, we confirm that Coralscapes is a challenging benchmark due to the imbalance in class occurrence and feature details. The Coralscapes dataset aims to catalyze the development of computer vision applications in coral reef science and conservation. Future work will extend Coralscapes to further increase the diversity (e.g. new biogeographic regions, different camera models, onboard lighting), and add more fine-grained classes, also including fish genera or species, and include instance-level annotations for fish.
VII Acknowledgments
We thank Dr. Ibrahim Souleiman Abdallah (University of Djibouti, Djibouti), Prof. Osama S. Saad, Mustafa Altaib Mohammed, and Abualdrdaa Mirgani Hajbakhit Omer (Red Sea University, Port Sudan), as well as Dr. Zekeria Zekeria and Temesgen Gebremeskel (Mai-Nefhi College of Asmara, Eritrea) for helping to acquire video material. We thank Dr. Ali Al-Sawalmih and Tariq Al-Salman (Marine Science Station, Aqaba), Prof. Maoz Fine, Nahum Sela (InterUniversity Institute of Marine Science, Eilat), Dr. Assaf Zvuloni (Nature Reserve Authority Israel), the Aqaba Special Economic Zone Authority of Jordan, the Ministry of Environment and Sustainable Development of Djibouti and the University of Djibouti, the Ministry of Marine Resources of Eritrea and the Mai-Nefhi College of Asmara, Eritrea for their support for enabling us to collect the videos. We thank Chiara Freneix, Géza Soldati, Camille Perrin, Amélie Menoud, Ines Stiti, Martin Métier, and Antoine Carnal for help with the frame annotations. The main body of data in this study were collected in the framework of expeditions organized by Transnational Red Sea Center, which is hosted by the Laboratory for Biological Geochemistry at EPFL. This work was funded in part by FNS grant 205321_212614, as well as EPFL and the Transnational Red Sea Center.
Appendix A Appendix: Implementation Details
A-A Benchmarking
@@
All models are implemented in PyTorch, with convolutional architectures (UNet++, DeepLabV3+) built using the segmentation-models-pytorch library [37] and transformer-based architectures (SegFormer, DPT) implemented using the Hugging Face transformers library [80]. Parameter efficient fine-tuning with LoRA was done using the Hugging Face PEFT library [52]. Training and evaluation are conducted on a system with a h100 GPU. The code used to train and evaluate the models is available on GitHub.  [^3] https://github.com/eceo-epfl/coralscapesScripts
@@
For the SegFormer models, training is conducted following [81], using a batch size of 8 when using the MiT-B2 backbone and 4 when using the MiT-B5 backbone for 1000 epochs, using the AdamW optimizer with an initial learning rate of 6e-5 (multiplied by 10 when using LoRA), weight decay of 1e-2 and polynomial learning rate scheduler with a power of 1. During training, images are randomly scaled within a range of 1 and 2, flipped horizontally with a 0.5 probability and randomly cropped to 1024×1024 pixels. Input images are normalized using the ImageNet mean and standard deviation. For evaluation, a non-overlapping sliding window strategy is employed, using a window size of 1024x1024 and a stride of 1024. When performing parameter efficient fine-tuning with LoRA, a rank size and alpha of 128 was used. We tried using ranks of 64 and 256, however that had no significant effect on the results. When training on the train+validation set and evaluating on the test set, we trained the model for 265 epochs for the MiT-B2 backbone, 75 epochs for the MiT-B5 backbone and 765 epochs for both backbones trained with LoRA. To increase the augmentation strength, we change the crops of size 1024 to random resized crops of scale 1.02 to 2.0 with aspect ratio 3/4 to 4/3, include random rotations of up to 15 degrees (cropped to exclude non-image areas), as well as random color jitters of contrast, saturation, and brightness between 0.8 and 1.2, and hue changes between -0.05 and 0.05. This model is trained for 100 epochs.
@@
For the DPT models [66] with a DINOv2 backbone, we followed the training procedure inspired by the Linear DINOv2 approach described above and the implementation of DepthAnythingV2. As such the model is trained for 1000 epochs with a batch size of 16 when using the base backbone and 8 when using the giant backbone, using the AdamW optimizer with an initial learning rate of 5e-5 (5e-6 for the backbone layers; multiplied by 10 when using LoRA), weight decay of 1e-2, and a polynomial learning rate scheduler with a power of 1. During training, images are resized to 1036×518 pixels, randomly horizontally flipped with a 0.5 probability, randomly cropped to size 518x518 and normalized using the ImageNet mean and standard deviation. For validation and testing, images are resized to the same dimensions and normalized similarly, while keeping the ground truth masks at their original size, using the same performance evaluation procedure as for the Linear DINOv2 model. When training on the train+validation set and evaluating on the test set, we trained the model for 165 epochs for the DINOv2-Base backbone, 55 epochs for the DINOv2-Giant backbone and 365 epochs for both backbones trained with LoRA.
A-B UCSD Mosaics Transfer Learning
All models are trained using the same optimizer and learning rate setting as in the Coralscapes benchmarking experiments. During training, only random horizontal and vertical flips of the 512$\times$512px patches are used during training as data augmentations. The models are trained for 100 epochs in the 5 & 10 labels per image setting, and for 200 epochs otherwise. Example training labels in each of the settings are shown in Figure 9, and example model outputs for DeepLabV3+ trained in the sparsest setting are shown in Figure 10.
@@ A-B UCSD Mosaics Transfer Learning
A-C Crown-of-Thorns Starfish Survey
@@
Appendix B Additional Samples

### v2

@@ 6 Conclusion and Future Work
We present the Coralscapes dataset for semantic understanding in reef scenes, which is the first general purpose dataset of coral scenes that are densely and consistently annotated by trained experts. Coralscapes is representative of the challenges of semantic segmentation in reef scenes, including complex scenes with challenging environmental conditions. By benchmarking state-of-the-art computer vision models, we confirm that Coralscapes is a challenging benchmark due to the imbalance in class occurrence and feature details. The Coralscapes dataset aims to catalyze the development of computer vision applications in coral reef science and conservation. Future work will extend Coralscapes to further increase the diversity (e.g. new biogeographic regions, different camera models, onboard lighting), and add more fine-grained classes, also including fish genera or species, and include instance annotations for fish.
Appendix A Appendix: Additional Dataset Information
Here, we give further information about the Coralscapes dataset and the annotated classes. A map of all reefs from which annotated data was collected is shown in Fig. 6. All videos were collected in reefs that are shallower than 15m, with the vast majority of images taken at a depth of around 5-7m. The distance from the seabed varies between roughly 50cm and 5m, with the vast majority of images taken at a surface distance of 1-2 meters. The images include diverse scenes from healthy reefs (approx. 70% live coral cover) to devastated reefs with 0% live coral cover, and include varying turbidity conditions (visibility from approx. 20m to 1.5m). Imagery was collected at all times of the day between just after sunrise to just before sunset in sunny and cloudy conditions. No artificial illumination or physical red filter was included.
Figure 6: Map of all reefs from which data was collected and annotated as part of Coralscapes.
We show the number of annotated pixels per class in Fig. 7, and the image frequency of each classes (on what percentage of images this class is present) in Fig. 8. All main reef-building coral types found in the Red Sea are represented in Coralscapes. Even though all the classes of Coralscapes can be found across the entire Indo-Pacific, there are morphotypes, species or genera falling into each class that are not adequately represented within Coralscapes as they do not (or rarely) appear in the Red Sea (e.g. Pocillopora grandis within the ‘pocillopora’ class or Xestospongia testudinaria within ‘sponge’).
Figure 7: Number of annotated pixels by class on a log-scale (with linear proportions for train/val/test).
Figure 8: Percentage of images with each class present.
Choosing the best resolution for a semantic segmentation dataset is challenging, as the label set should capture all classes of interest while at the same time having enough samples of each class to be able to accurately assess the performance of a machine learning system. In Figure 9, we propose a hierarchical scheme for summarizing the label set into coarser classes.
Figure 9: Proposed label hierarchy for summarization.
A.1 Annotation Class Guide
Table Acropora Alive
X Acropora that grow in a tabular growth form.
Acropora Alive
X Acropora that do not grow in tabular growth form.
Stylophora Alive
X Clearly identifiable Stylophora.
Pocillopora Alive
X Clearly identifiable Pocillopora
Branching Alive
X Branching corals that can surely be determined to be alive, but do not fit in to ‘table acropora alive‘, ‘acropora‘, ‘stylophora’, or ‘pocillopora’ because they are from a different genus or appear slightly blurred in the context.
Turbinaria
X Colloqually the ‘scroll’ coral. Does not include the macroalgae genus Turbinaria.
Meandering Alive
X Corals of a meandering growth form. Includes Platygyra, Lobophyllia, Symphyllia.
Massive/Meandering Alive
X Corals in a massive growth form. Prominently includes Porites, Favia, Favites, and many others. Includes some likely meandering corals (like Platygyra) where the meandering structure can not be clearly identified.
Millepora
X The ‘fire coral’ Millepora is technically not a coral, but a hydrozoan. Appears most commonly in a branching form (Millepora Dichotoma).
Other Coral Alive
X All other live corals. Includes corals in thin plate or encrusting growth form, soft corals, and corals that can not be clearly classified into the other classes.
Table Acropora Dead
X Dead Acropora tables which are dead or collapsed but are not yet overgrown by algae. Parts visibly overgrown by algae are labeled as ‘algae covered substrate’.
Branching Dead
X Other dead branching coral, including acropora, pocillopora, and stylophora. When overgrown by algae, labeled as ‘algae covered substrate’.
Meandering Dead
X Dead meandering corals (Lobophyllia, Symphyllia, Platygyra, etc.).
Massive/Meandering Dead
X Dead massive corals or meandering corals that have decomposed enough so that the meandering structure is no longer visible.
Other Coral Dead
X All other dead coral skeletons.
Branching Bleached
X Bleached branching coral of all kinds, including bleached table acropora.
Meandering Bleached
X Bleached meandering corals (Lobophyllia, Symphyllia, Platygyra, etc.).
Massive/Meandering Bleached
X Bleached massive corals or meandering corals that have decomposed enough so that the meandering structure is no longer visible.
Other Coral Bleached
X All other bleached coral, including bleached soft coral.
Sponge
X Sponges of all kind.
Crown of Thorns
X Acanthaster planci, known to cause outbreaks in which it can severely reduce coral cover.
Sea Urchin
X
Sea Cucumber
X
Anemone
X
Clam
X Live giant clams.
Dead Clam
X Dead giant clams.
Other Animal
X Includes starfish (except the crown-of-thorns starfish), feather worms, sea turtles, and other non-identifiable invertebrates or animals of which there are not enough annotations to warrant a separate class.
Trash
X Includes all kinds of marine litter. Most common are plastic items including bags, cups, and bottles, aluminum cans and glass bottles, as well as abandoned fishing material, and parts of boats and machines.
Transect Line
X Rolled out transect tape. Excludes reels or other strands of rope that are laid out.
Transect Tools
X This includes transect reels & spools, tags and markers placed on the reef, diving weights, surface marker buoys and the string attaching them to the ground or weights on the ground.
Human
X The human class includes divers and snorkelers, of which sometimes only a hand or a fin is in the frame. Some ambiguity can arise when a human carries a transect tool that is not laid out on the ground, like a transect reel. In these cases, we generally decide this tool becomes part of the human polygon.
Fish
X Fish of all kinds.
Background
X This class is assigned to pixels that are too far away or too blurry to be classifiable into any other class. This includes the water surface, which is sometimes visible. If Coralscapes is used in the context of benthic mapping, we recommend masking or excluding background.
Dark
X Parts of the image that are too dark to discern the benthic class, but that are within the substrate. Includes holes, crevices, cracks, or dark shadows cast onto the substrate. If Coralscapes is used in the context of benthic mapping, dark areas should be included in the outputs.
Sand
X Loose, fine sand.
Rubble
X Small loose fragments of rocky substrate or dead coral.
Algae Covered Substrate
X Substrate covered in turf algae or other macroalgae, including fleshy algae and Turbinaria.
Hard Substrate
X Hard substrate that is part of the reef, which can not be identified into any of the other classes. Includes rocks and heavily decomposed coral skeletons. Also includes human-made structures (underwater infrastructure) such as pipes, pier columns, coral nursery tables, buoys and their lines, as well as boat anchors.
Seagrass
X
Appendix B Appendix: Implementation Details
B.1 Benchmarking
@@
All models are implemented in PyTorch, with convolutional architectures (UNet++, DeepLabV3+) built using the segmentation-models-pytorch library [39] and transformer-based architectures (SegFormer, DPT) implemented using the Hugging Face transformers library [86]. Parameter efficient fine-tuning with LoRA was done using the Hugging Face PEFT library [55]. Training and evaluation are conducted on a system with a h100 GPU. The code used to train and evaluate the models is available on GitHub [^3] https://github.com/eceo-epfl/coralscapesScripts.
@@
For the SegFormer models, training is conducted following [87], using a batch size of 8 when using the MiT-b2 backbone and 4 when using the MiT-b5 backbone for 1000 epochs, using the AdamW optimizer with an initial learning rate of 6e-5 (multiplied by 10 when using LoRA), weight decay of 1e-2 and polynomial learning rate scheduler with a power of 1. During training, images are randomly scaled within a range of 1 and 2, flipped horizontally with a 0.5 probability and randomly cropped to 1024×1024 pixels. Input images are normalized using the ImageNet mean and standard deviation. For evaluation, a non-overlapping sliding window strategy is employed, using a window size of 1024x1024 and a stride of 1024. When performing parameter efficient fine-tuning with LoRA, a rank size and alpha of 128 was used. We tried using ranks of 64 and 256, however that had no significant effect on the results. When training on the train+validation set and evaluating on the test set, we trained the model for 265 epochs for the MiT-b2 backbone, 75 epochs for the MiT-b5 backbone and 765 epochs for both backbones trained with LoRA.
@@
For the DPT models [72] with a DINOv2 backbone, we followed the training procedure inspired by the Linear DINOv2 approach described above and the implementation of DepthAnythingV2 [88]. As such the model is trained for 1000 epochs with a batch size of 16 when using the base backbone and 8 when using the giant backbone, using the AdamW optimizer with an initial learning rate of 5e-5 (5e-6 for the backbone layers; multiplied by 10 when using LoRA), weight decay of 1e-2, and a polynomial learning rate scheduler with a power of 1. During training, images are resized to 1036×518 pixels, randomly horizontally flipped with a 0.5 probability, randomly cropped to size 518x518 and normalized using the ImageNet mean and standard deviation. For validation and testing, images are resized to the same dimensions and normalized similarly, while keeping the ground truth masks at their original size, using the same performance evaluation procedure as for the Linear DINOv2 model. When training on the train+validation set and evaluating on the test set, we trained the model for 165 epochs for the DINOv2-Base backbone, 55 epochs for the DINOv2-Giant backbone and 365 epochs for both backbones trained with LoRA.
B.2 UCSD Mosaics Transfer Learning
All models are trained using the same optimizer and learning rate setting as in the Coralscapes benchmarking experiments. During training, only random horizontal and vertical flips of the $\times512$px patches during training as data augmentations. The models are trained for 100 epochs in the 5 & 10 labels per image setting, and for 200 epochs otherwise. Example training labels in each of the settings are shown in Figure 10, and example model outputs for DeepLabV3+ trained in the sparsest setting are shown in Figure 11.
@@ B.2 UCSD Mosaics Transfer Learning
B.3 Crown-of-Thorns Starfish Survey
@@
Appendix C Comparison to CoralSCOP
In this Section we emphasize the key differences from Coralscapes to the dataset from CoralSCOP [91]:
  1. 1.
    General purpose: CoralSCOP only has segmentation masks for visible corals, and no other classes. This makes CoralSCOP unsuitable for many applications (e.g. marine robotics), where it is critical to segment other classes (divers, fish, other substrates) to avoid collisions and ensure safe deployment. Coralscapes goes beyond coral segmentation, and aims for scene understanding in coral reefs.
  2. 2.
    Data quality: Our data was collected as part of research & monitoring campaigns by the authors. We can attest that all images are original and authentic (not altered or generated by AI), whereas the majority of CoralSCOP is crawled from the internet or recompiled from existing sources.
  3. 3.
    Annotation quality: The Coralscapes segmentations follow a stricter quality standard than existing datasets w.r.t to coverage and a conservatively applied class hierarchy.
These key points position Coralscapes as a high-quality dataset for general purpose reef segmentation, whereas CoralSCOP’s main aim is breadth and coverage.
To supplement the findings of Fig. 3, which show that Coralscapes is more complex than CoralSCOP, we offer a direct comparison to CoralSCOP images that have a high proportion of the images annotated in Fig. 13. Of the first 105 CoralSCOP images with geqgeq90% cover (below), 98 are extreme close-ups of corals, and the remaining ones have severely imprecise segmentation masks: there is no clear distinction between dead and live coral, and often there are rocks & large dark patches which are mistakenly included in the coral mask. The images shown are more than 6% of CoralSCOP images with $\geq$90% annotation cover.
Figure 13: Samples from CoralSCOP. Top: 100 samples of images with $\geq$90% annotation cover, showing that the vast majority are extreme close-ups of corals. Bottom: for images that are not close-ups, the coral segmentation masks are coarse.
Appendix D Additional Samples